using the integral calculus to solve a simple engineering question,
a missile is accelerating at a rate of 6 t m/sec^2 from a position at rest in a silo 40m above ground level. how high above the ground will it be after 4 seconds

Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)
Answer by Alan3354(69443)

a missile is accelerating at a rate of 6t m/sec^2 from a position at rest in a silo 40m above ground level. how high above the ground will it be after 4 seconds
-----------------
40 meters above ground? Most are below ground.
-------------
Accel = 1st derivative of speed = 2nd derivative of displacement.
a = 6t (assuming t is time in seconds)
Using its starting height as zero at t = 0:
---
INT(6t) = 3t^2 + C
C = 0
-------
INT(3t^2) = t^3 + C
Again, C = 0
----------------
s = t^3
At t = 4, s = 64 meters
Add or subtract the 40 meters